1. The instructions for a solution ask you to put 15.0 grams of HCl (strong) into 2.0L of water solution, what is the expected pH of the solution?
1
Expert's answer
2020-06-04T10:37:35-0400

The molarity becomes;

"Molarity= mass\/rfm"

Molarity=15/36.46=0.41 M

"PH=-log(0.41)"

PH=3.8